# VCR: Visual Caption Restoration

### v1

@@ 4.4 Human Evaluation
We recruited 7 volunteers to conduct human evaluation on a subset of the samples of our datasets. Two out of the seven evaluators are native English speakers, while five are native Chinese speakers who are also fluent in English171717The participants’ TOEFL scores range from 102/120 to 112/120.. All the volunteers have earned post-graduate degrees majoring in one of biology, statistics, computer science, economy, etc. The evaluations were conducted on a voluntary basis and participants received no rewards.
@@ 5 Related Work
Masked language modeling (MLM) introduced by BERT [13] and its autoregressive counterparts by GPT [8] have been the foundations of pre-training modern natural language processing (NLP) models. Both are trained to predict a masked-out proportion based on the context. Similarly, inspired by BERT and GPT, iGPT [9] uses the input sequences of pixels to predict the next unknown pixels. Masked Autoencoders (MAE) [19] introduce Masked Image Modeling (MIM) by masking out random patches from input images and reconstructing the missing parts in the images in pixel space. The goal is to encourage the model to learn the representations of the image by understanding the context and structure of the visible patches, which helps reconstruct the masked images accurately.
@@
Vision-language models are designed for tasks that involve understanding and generating content from images and text [56, 39, 31, 32]. For example, models have been developed to combine Llama3 with advanced vision-language processing capabilities to handle complex multimodal tasks [70, 66, 24, 69, 63, 14]. Qwen-VL [4] enhances visual-linguistic representations for more accurate contextual interpretations, while OpenGVLab-InternVL-Chat [11, 10] merges the InternVL framework with interactive chat capabilities. General-purpose models like GPT-4 series models [50, 49], Claude series models [2] have also been adapted for vision-language tasks, demonstrating strong performance in multimodal tasks. Finally, DocLLM [60] specializes in document understanding by integrating visual and textual data to enhance the interpretation and generation of document-related content. These models collectively represent significant advancements in vision-language integration, contributing unique capabilities and enhancements to the understanding and generation of multimodal information.
@@ 6 Conclusion
We conducted extensive evaluations of state-of-the-art vision-language models (VLMs) in both English and Chinese. The results demonstrated significant room for improvement, suggesting that current models have not yet fully harnessed the capabilities necessary for VCR. We selected models representing both the highest and average performance tiers for additional fine-tuning with our dataset. Although fine-tuning exhibited potential for enhancing VCR capabilities, it did not consistently result in significant improvements, indicating the complexity and challenges of adapting models to this task.
By introducing the VCR task and its specialized dataset, we aim to advance research in vision-language interaction. The unique challenges of VCR seek to enhance model development and training, extending the limits of multimodal AI. We invite the community to utilize our dataset and develop innovative strategies to boost VLM model performance.
@@ Appendix A Additional evaluation results on first 100 and 500 test cases
EN Easy Closed-source models
@@
ZH Easy Closed-source models
@@
EN Easy Closed-source models
@@
ZH Easy Closed-source models

### v2

@@ 4.4 Human Evaluation
We recruited 7 volunteers to perform human evaluation on a subset of the samples of our datasets. Two out of the seven evaluators are native English speakers, while five are native Chinese speakers who are also fluent in English171717The TOEFL scores of the non-native English-speaking participants range from 102/120 to 112/120.. All volunteers have earned postgraduate degrees majoring in one of the following fields: biology, statistics, computer science, and economics. The evaluations were conducted on a voluntary basis and participants received no rewards.
@@ 5 Related Work
Masked language modeling (MLM) introduced by BERT [13] and its autoregressive counterparts by GPT [8] have been the foundations of pre-training modern natural language processing (NLP) models. Both are trained to predict a masked-out proportion based on the context. Similarly, inspired by BERT and GPT, iGPT [9] uses the input sequences of the pixels to predict the next unknown pixels. Masked Autoencoders (MAE) [19] introduce Masked Image Modeling (MIM) by masking out random patches from input images and reconstructing the missing parts in the images in pixel space. The goal is to encourage the model to learn the representations of the image by understanding the context and structure of the visible patches, which helps to reconstruct the masked images accurately.
@@
Vision-language models are designed for tasks that involve understanding and generating content from images and text [58, 40, 31, 32]. For example, models have been developed to combine Llama3 with advanced vision-language processing capabilities to handle complex multimodal tasks [75, 71, 24, 74, 67, 14]. Qwen-VL [4] enhances visual-linguistic representations for more accurate contextual interpretations, while OpenGVLab-InternVL-Chat [11, 10] merges the InternVL framework with interactive chat capabilities. These studies typically employ a multimodal encoder [53, 76, 70] to process multimodal data, which is then mapped to the same input space of the language model. General-purpose models such as the GPT-4 series models [51, 50], the Claude series models [2], the Gemini series models [59] and the Reka series models [60] have also been adapted for vision-language tasks, demonstrating strong performance in multimodal tasks. Finally, DocLLM [64] specializes in document understanding by integrating visual and textual data to enhance the interpretation and generation of document-related content. These models collectively represent significant advancements in vision-language integration, contributing unique capabilities and enhancements to the understanding and generation of multimodal information.
@@ 6 Conclusion
We conducted extensive evaluations of state-of-the-art vision-language models (VLMs) in both English and Chinese. The results demonstrated significant room for improvement, suggesting that current models have not yet fully exploited the capabilities necessary for VCR. We selected models representing both the highest and average performance tiers for additional fine-tuning with our dataset. Although fine-tuning exhibited potential for enhancing VCR capabilities, it did not consistently result in significant improvements, indicating the complexity and challenges of adapting models to this task.
By introducing the VCR task and its specialized dataset, we aim to advance research in vision-language interaction. The unique challenges of VCR seek to improve model development and training, extending the limits of multimodal AI. We invite the community to utilize our dataset and develop innovative strategies to boost the performance of vision-language models.
@@ Appendix A Additional evaluation results on first 100 and 500 test cases
English Easy Closed-source models
@@
Chinese Easy Closed-source models
Claude 3 Opus 0 2.2
Claude 3.5 Sonnet - -0.53 -1.84
Gemini 1.5 Pro - 0.53 0.16
GPT-4o - -6.91 -9.72
GPT-4 Turbo - 0.53 3.58
@@
Claude 3 Opus - 0.53 1.45
Claude 3.5 Sonnet - 0.53 0.79
Gemini 1.5 Pro - 0 -1.76
GPT-4o - 1.06 0
GPT-4 Turbo - -0.53 0.49
@@
English Easy Closed-source models
@@
Chinese Easy Closed-source models
Claude 3 Opus - -0.1 1.49
Claude 3.5 Sonnet - 0.2 0.03
Gemini 1.5 Pro - 0.6 -0.37
GPT-4o - -7.58 -9.19
GPT-4 Turbo - 0.1 1.45
@@
Claude 3 Opus - 0.2 1.13
Claude 3.5 Sonnet - 0.2 1.63
Gemini 1.5 Pro - 0.2 0.07
GPT-4o - 0.4 -0.17
GPT-4 Turbo - -0.2 1.72